# UniTalker: Conversational Speech-Visual Synthesis

### v1

@@ 6.2. Synthesized Speech Evaluation
| -w/o E-guided | 0.894 | 50.367 | 0.682 | 3.925±0.026 | 3.916±0.015 | 26.376 | 0.753 | 6.284 | 4.015±0.037 | 4.026±0.016 | 4.008±0.021 |
| -w/o | 0.871 | 54.964 | 0.642 | 3.946±0.031 | 3.894±0.016 | 28.657 | 0.748 | 6.103 | 3.987±0.036 | 3.994±0.018 | 3.975±0.019 |
| -w/o | 0.826 | 65.845 | 0.597 | 3.672±0.043 | 3.608±0.017 | 60.486 | 0.607 | 5.163 | 3.707±0.018 | 3.709±0.043 | 3.768±0.016 |
| -w/o Stage-2 | 0.751 | 66.238 | 0.601 | 3.601±0.026 | 3.643±0.029 | 50.781 | 0.657 | 5.667 | 3.735±0.028 | 3.807±0.013 | 3.792±0.037 |
| -w/o Stage-3 | 0.886 | 45.238 | 0.733 | 4.015±0.017 | 4.083±0.024 | 24.961 | 0.796 | 6.194 | 4.093±0.016 | 4.116±0.009 | 4.103±0.028 |
| -w/o Alignment | 0.781 | 72.684 | 0.613 | 3.693±0.022 | 3.683±0.032 | 59.127 | 0.572 | 4.452 | 3.567±0.019 | 3.659±0.021 | 3.612±0.016 |
| -w/o D-history | 0.853 | 60.123 | 0.587 | 3.737±0.014 | 3.642±0.026 | 53.127 | 0.648 | 6.127 | 3.867±0.034 | 3.827±0.013 | 3.646±0.022 |
| -w/o U | 0.892 | 44.651 | 0.738 | 4.103±0.011 | 4.006±0.028 | 25.168 | 0.801 | 6.343 | 4.112±0.013 | 4.032±0.025 | 4.167±0.023 |
| -w/o U | 0.862 | 48.367 | 0.712 | 3.989±0.037 | 3.972±0.015 | 26.367 | 0.785 | 6.313 | 4.098±0.022 | 3.976±0.018 | 3.986±0.037 |
| -w/o U | 0.889 | 43.194 | 0.732 | 4.068±0.016 | 4.027±0.041 | 28.319 | 0.762 | 6.178 | 3.924±0.014 | 3.984±0.031 | 3.994±0.035 |
| -w/o U | 0.896 | 44.128 | 0.718 | 4.038±0.019 | 4.037±0.006 | 25.367 | 0.768 | 6.245 | 4.035±0.016 | 4.086±0.006 | 4.102±0.013 |
| -w/o P1→N-1 | 0.871 | 45.328 | 0.726 | 4.062±0.023 | 4.052±0.011 | 24.869 | 0.798 | 6.271 | 4.007±0.022 | 4.052±0.014 | 4.156±0.026 |

### v2

@@ 6.2. Synthesized Speech Evaluation
-w/o E-guided | 0.894 | 50.367 | 0.682 | 3.925±0.026 | 3.916±0.015 | 26.376 | 0.753 | 6.284 | 4.015±0.037 | 4.026±0.016 | 4.008±0.021 |
-w/o | 0.871 | 54.964 | 0.642 | 3.946±0.031 | 3.894±0.016 | 28.657 | 0.748 | 6.103 | 3.987±0.036 | 3.994±0.018 | 3.975±0.019 |
-w/o | 0.826 | 65.845 | 0.597 | 3.672±0.043 | 3.608±0.017 | 60.486 | 0.607 | 5.163 | 3.707±0.018 | 3.709±0.043 | 3.768±0.016 |
-w/o Stage-2 | 0.751 | 66.238 | 0.601 | 3.601±0.026 | 3.643±0.029 | 50.781 | 0.657 | 5.667 | 3.735±0.028 | 3.807±0.013 | 3.792±0.037 |
-w/o Stage-3 | 0.886 | 45.238 | 0.733 | 4.015±0.017 | 4.083±0.024 | 24.961 | 0.796 | 6.194 | 4.093±0.016 | 4.116±0.009 | 4.103±0.028 |
-w/o Alignment | 0.781 | 72.684 | 0.613 | 3.693±0.022 | 3.683±0.032 | 59.127 | 0.572 | 4.452 | 3.567±0.019 | 3.659±0.021 | 3.612±0.016 |
-w/o D-history | 0.853 | 60.123 | 0.587 | 3.737±0.014 | 3.642±0.026 | 53.127 | 0.648 | 6.127 | 3.867±0.034 | 3.827±0.013 | 3.646±0.022 |
-w/o U | 0.892 | 44.651 | 0.738 | 4.103±0.011 | 4.006±0.028 | 25.168 | 0.801 | 6.343 | 4.112±0.013 | 4.032±0.025 | 4.167±0.023 |
-w/o U | 0.862 | 48.367 | 0.712 | 3.989±0.037 | 3.972±0.015 | 26.367 | 0.785 | 6.313 | 4.098±0.022 | 3.976±0.018 | 3.986±0.037 |
-w/o U | 0.889 | 43.194 | 0.732 | 4.068±0.016 | 4.027±0.041 | 28.319 | 0.762 | 6.178 | 3.924±0.014 | 3.984±0.031 | 3.994±0.035 |
-w/o U | 0.896 | 44.128 | 0.718 | 4.038±0.019 | 4.037±0.006 | 25.367 | 0.768 | 6.245 | 4.035±0.016 | 4.086±0.006 | 4.102±0.013 |
-w/o P1→N-1 | 0.871 | 45.328 | 0.726 | 4.062±0.023 | 4.052±0.011 | 24.869 | 0.798 | 6.271 | 4.007±0.022 | 4.052±0.014 | 4.156±0.026 |